Logistics Solutions provides order fulfillment services for dot.com merchants. The company maintains warehouses that stock items carried by its dot.com clients. When a client receives an order from a customer, the order is forwarded to Logistics Solutions, which pulls the item from storage, packs it, and ships it to the customer. The company uses a predetermined variable overhead rate based on direct labor-hours.

In the most recent month, 190,000 items were shipped to customers using 8,300 direct labor-hours. The company incurred a total of $29,050 in variable overhead costs.

According to the company’s standards, 0.03 direct labor-hours are required to fulfill an order for one item and the variable overhead rate is $3.55 per direct labor-hour.

Required:

1. What is the standard labor-hours allowed (SH) to ship 190,000 items to customers?

2. What is the standard variable overhead cost allowed (SH × SR) to ship 190,000 items to customers?

3. What is the variable overhead spending variance?

4. What is the variable overhead rate variance and the variable overhead efficiency variance?
Answer #1

(1) standard hours allowed for 190000 items

=190000 items*0.03 direct labor hours per item

=5700 hiours

(2)SH * SR

=5700*3.55$

=20235$

(3)variable overhead spending variance

Actual hours*actual rate - standard hour*standard rate

=29050-20235

=8815$ unfavorable [ as the actual overhead is more than standard the variance is unfavorable]

(4)variable rate variance

(actual rate-standard rate)*Actual quantity

29050-3.55*8300

=29050-29465

=415$ favorable as actual per hour rate is lesser than standard rate

variable overhead efficiency variance

Actual hour*standard rate -standard hour*standard rate

8300*3.55-5700*3.55

=29465-20235

=9230$ unfavorable